House A is three times as old as house b. six years ago house a was four times as old as house b. how old is each house
Answer by sachi(548)

House A is three times as old as house b.
let house b is x yrs old then house a is 3x yrs old
six years ago house a was four times as old as house b.
so 3x-6=4(x-6)=4x-24
or x=24-6=18
so house b is 18 yrs old then house a is 54 yrs old
ans